OpenStudy (anonymous):

describe a pattern in the sequence of numbers. what is the next number 7, 9, 13, 19, 27

OpenStudy (anonymous):

each successive element is 2n + previous element, where n = 1,2,3,4,5,.....

jimthompson5910 (jim_thompson5910):

First time you're adding 2 (7+2 = 9) Then you're adding 4 (9+4 = 13) Then you're adding 6 (13+6 = 19) Then you're adding 8 (19+8 = 27) See the pattern?

OpenStudy (jamesj):

Seriously? What do you think the answer is here?

OpenStudy (anonymous):

Add 10 to the last one?

OpenStudy (anonymous):

i got 37

jimthompson5910 (jim_thompson5910):

you got it, you add 10 to 27 to get 37

OpenStudy (anonymous):

yes i put add by even numberr